OpenStudy (mackenzie2013):

Can someone check this ? Find the area of a circle that has a radius of 14 feet. Round your answer to the nearest hundredth. A=pi R^2 3.14x14=43.96 Then what?

OpenStudy (anonymous):

You forgot one step. Before multiply 3.14 with the radius you have to square it

OpenStudy (anonymous):

@Mackenzie2013 look at the equation. u wrote A=pi time r^2 but u didn't do it. 14^2=196....196 times pi=..?

OpenStudy (anonymous):

Well square the radius, not pi

OpenStudy (mackenzie2013):

Wait! I seee!

OpenStudy (mackenzie2013):

So then you take 196 and times it by 3.14

OpenStudy (anonymous):

yes

OpenStudy (mackenzie2013):

@lacrosseplayer22 can you help me out with one more?

OpenStudy (anonymous):

sure

OpenStudy (mackenzie2013):

Find the area of a circle whose circumference is 31.4 centimeters. A =

OpenStudy (anonymous):

The circumference of a circle is C=2pi(r) we are given that 31.4=2pi(r) to find the radius we must divide 31.4 by 2pi, can you do that?

OpenStudy (anonymous):

yes

OpenStudy (mackenzie2013):

2pi is that pi times two?

OpenStudy (anonymous):

yes it is

OpenStudy (mackenzie2013):

0.2?

OpenStudy (anonymous):

so if radius is 5 then the are should be 78.53.... right @Venny

OpenStudy (anonymous):

I am here calm down a second Yes the radius is 5, area of a circl A=pi(r)^2 A=pi(5)^2 A=25pi A=78.54

OpenStudy (anonymous):

sorry

OpenStudy (mackenzie2013):

So 78.54 is the answer?

OpenStudy (anonymous):

Yes that is, also 0.2 is not the same thing as *2, 0.2 is the same thing as 2/10 which is less than 1 so it makes it smaller.

OpenStudy (mackenzie2013):

It says its wrong..

OpenStudy (anonymous):

?? that doesn't make sense

OpenStudy (anonymous):

Yeah we didn't mess up anywhere

OpenStudy (mackenzie2013):

Somethings gotta be messed up.

OpenStudy (mackenzie2013):

Find the area of a circle whose circumference is 31.4 centimeters. A = cm2

OpenStudy (mackenzie2013):

Thats the question

OpenStudy (anonymous):

Did you forget to label it with cm^2?

OpenStudy (mackenzie2013):

Yes ... Facepalm

OpenStudy (mackenzie2013):

wait no

OpenStudy (mackenzie2013):

okay this is how it was typed. Find the area of a circle whose circumference is 31.4 centimeters. A = cm2

OpenStudy (mackenzie2013):

got it!

OpenStudy (anonymous):

oh no she's right. im wrong. if circumference=pi times diameter and Area= Pi times r^2 then we do 31.4 divided by 3.14 is 10. so we know r^2 is 10. to find r we have to find the root of 10. then we do A=pi r ^2

OpenStudy (anonymous):

face palm goes to me

OpenStudy (anonymous):

2r=diameter